OpenStudy (anonymous):

3x + 2y = 4 2x - y = 5 This system of equations A)has no solution. B)has one solution. C)is coincident.

OpenStudy (amoodarya):

multiply second by 2 then add them emit y so you will find x then put x into one of the equation so you will have value of y

OpenStudy (amoodarya):

3x+2y=4 2x-y=5 ----> 4x-2y=10 adding 3x+2y+4x-2y=4+10 7x=14 x=2 now put x=2 in one of them 3x+2y=4 3(4)+2y=4 2y=4-12 2y=-8 y=-4

OpenStudy (anonymous):

so it has one solution right?

OpenStudy (nirmalnema):

you can also check it by just comparing coefficients for eg... ax+by+c=0 ......(1) dx+ey+f=0....(2) if a/d is not equal to b/e then equations has only one solution if a/d is equal to b/e which is not equal to c/f the equations has no solutions if all are equal in above case the it is coincident